Dozier Company produced and sold 1,000 units during its first month of operations. It reported the following costs and expenses for the month: $ 83,000 S 42,000 Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Total manufacturing overhead Variable selling expense Fixed selling expense Total selling expense Variable administrative expense Fixed administrative expense Total administrative expense $ 20,600 32,200 $ 52,800 S 14, 800 23, 600 s 38,400 S 5, 400 27,800 $ 33, 200 Required 1. With respect to cost classifications for preparing financial statements a. What is the total product cost? b. What is the total period cost? 2. With respect to cost classifications for assigning costs to cost objects: a. What is total direct manufacturing cost? b. What is the total indirect manufacturing cost? 3. With respect to cost classifications for manufacturers a. What is the total manufacturing cost? b. What is the total nonmanufacturing cost? C. what is the total conversion cost and prime cost? 4. With respect to cost classifications for predicting cost behavior a. What is the total variable manufacturing cost? b. What is the total fixed cost for the company as a whole? c. What is the variable cost per unit produced and sold? 5. With respect to cost classifications for decision making: a. If Dozier had produced 1,001 units instead of 1,000 units, how much incremental manufacturing cost would it have incurred to make the additional unit?
Answer #1
Solution:
1a. Total product cost 177,800
1b. Total period cost 71,600
Working Notes:
Total product cost
Direct materials 83,000 A
Direct labor 42,000 B
Total manufacturing overhead 52,800 C
Total product cost 177,800 D=A+B+C
Total period cost
Total selling expense 38,400 A
Total administrative expense 33,200 B
Total period cost 71,600 C =A + B
2a. Total Direct manufacturing cost 125,000
2b. Total Indirect manufacturing cost 52,800
Working Notes:
Total Direct manufacturing cost
Direct materials 83,000 A
Direct labor 42,000 B
Total Direct manufacturing cost 125,000 C=A+B
Total Indirect manufacturing cost
Variable manufacturing overhead 20,600 A
Fixed manufacturing overhead 32,200 B
Total Indirect manufacturing cost 52,800 C=A+B
3a. Total manufacturing cost 177,800
3b. Total nonmanufacturing cost 71,600
3c. Total conversion cost 94,800
Total prime cost 125,000
Working Notes:
Total manufacturing cost
Direct materials 83,000 A
Direct labor 42,000 B
Total manufacturing overhead 52,800 C
Total manufacturing cost 177,800 D=A+B+C
Total non manufacturing cost
Total selling expense 38,400 A
Total administrative expense 33,200 B
Total nonmanufacturing cost 71,600 C =A + B
Total conversion cost
Direct labor 42,000 A
Total manufacturing overhead 52,800 B
Total conversion cost 94,800 C =A + B
Total prime cost
Direct materials 83,000 A
Direct labor 42,000 B
Total prime cost 125,000 C =A + B
4a. Total variable manufacturing cost 145,600
4b. Total fixed cost 83,600
4c. Variable cost per unit produced and sold 166
Working Notes:
Total variable manufacturing cost
Direct materials 83,000 A
Direct labor 42,000 B
Variable manufacturing overhead 20,600 C
Total variable manufacturing cost 145,600 D=A+B+C
Total fixed cost
Fixed manufacturing overhead 32,200 A
Fixed selling expense 23,600 B
Fixed administrative expense 27,800 C
Total fixed cost 83,600 D=A+B+C
Variable cost per unit produced and sold
Direct materials 83,000 A
Direct labor 42,000 B
Variable manufacturing overhead 20,600 C
Variable selling expense 14,800 D
Variable administrative expense 5,400 E
Total variable cost 165,800 F=A+B+C+D+E
Number of units produced and sold 1000 G
Variable cost per unit produced and sold 165.8 H = F ÷ G
166
5a. Incremental cost per unit produced 146
Working Notes:
Direct materials 83,000 A
Direct labor 42,000 B
Variable manufacturing overhead 20,600 C
Total incremental cost 145,600 D=A+B+C
Number of units produced and sold 1,000 E
Incremental cost per unit produced 145.6 F = D ÷ E
So for additional 1 unit cost per unit 146
Please feel free to ask if anything about above solution in comment section of the question.